# Unbundle-Rewrite-Rebundle: Runtime Detection and Rewriting of Privacy-Harming Code in JavaScript Bundles

### v2

@@ 3.1.1. Gathering JavaScript resources
In order to create a valid parser for bundles in the wild, we gathered examples of resources loaded in popular websites. We developed a puppeteer-based crawler that, upon visiting a page, intercepts network requests and stores a copy of observed script responses. We visited domains from the Tranco list (Le Pochat et al., 2019) and gathered 30,930 scripts from 1,063 sites (1K crawl).
@@ 3.1.2. Generating ASTs
For each of the gathered scripts, URR uses acorn, a community-developed, open-source JavaScript parser to generate an AST (acornjs, 2023). The generated AST complies with the ESTree Spec to ensure a consistent, standardized representation that can be reproduced by alternative implementations (McKenzie et al., 2023).
@@
We began with an understanding of the general structure of the output of a webpack bundle (see Listing 1). We developed a script that parses an AST and looks for a module array or object, i.e., an array or object that comprises functions. We used this initial logic and adopted a code optimization methodology based on ground-truth gathered through expert manual annotation. We describe the process below.
@@
    We manually evaluated the plaintext script and the AST of each sampled resource. In doing so, we checked for the presence of webpack-specific code patterns. We annotated each resource with a boolean value indicating whether we determined the script as a webpack bundle (Annotation 1). The manual checks included:
@@
    We used the refined version of our code to annotate the same sample and noted any improvements. This refined version of the code as the initial version of the subsequent phase, against new samples.
@@
We used the resulting logic to annotate all scripts gathered in our 1K crawl. We annotated a total of 11,995 scripts as webpack bundles. Finally, we performed a spot check of the generated annotation. We randomly sampled and manually annotated 300, previously unsampled scripts of which we identified 99 scripts as webpack bundles. Upon comparing and verifying the code-assigned annotation, we observed a 97% precision rate in identifying webpack bundles and a 99.8% precision rate in gathering component modules.
@@ 3.2. Processing Modules
Each module gathered from the previous phase comprises a sub-tree, i.e., a partial AST of the larger AST representation of the script. The module’s AST contains information about both, the structure of the underlying code and associated names of variables, functions, and properties as included within the script. Script attributes like variable names are extremely volatile and have limited use in reliably identifying target modules. URR therefore only considers (1) the structure of the AST, i.e., the parent-child relationships between the nodes that comprise the AST, and (2) the type of each AST node, i.e., attributes which comply with the ESTree Spec, and are hence limited to a deterministic set of values.
To distill and represent only these specific attributes from a module’s AST, we adopt a version of a cryptographic concept used in integrity verification: Merkle trees (Merkle, 1987). As an example, consider a list of data blocks that need to be verified (e.g., transactions in a blockchain). We represent this list of blocks in a tree structure. Consider a tree where each leaf node represents a piece of data, and each non-leaf node is a cryptographic hash of its child nodes. These hashes propagate upwards, converging into a singular root hash, known as the Merkle root. Any alteration in the foundational data triggers a modification in the root hash, instantly signaling tampering or modifications. However, if the root hash is verified, all of its children are verified as well.
@@
    For every child node, it recursively calls the function to gather the child node’s hash. The hash of each child in-turn results from the hashes of its children.
@@
    To generate a hashed representation of the root node, URR concatenates the type of the node to the sum of hashes of its children, and hashes the resulting string.
@@
This processed representation now comprises an alternative tree representation with equal depth to its original counterpart. However, tree-based comparisons are complex and add a large performance overhead. Recall that while processing an AST, top-level nodes contain information about underlying children. Therefore, comparing a limiting comparison to a few high-level nodes of the processed module can provide insight into the similarity of the module’s AST. To this end, URR gathers a list of top-level nodes, i.e., given a module with webpack’s function wrappers, it gathers a list of root nodes of each subtree corresponding to a top-level statement within the function. Additionally, URR associates each entry in this list with a weight that represents the number of child nodes it represents, i.e., it weighs a node representing a complex function higher than a node representing a simple variable declaration.
@@ 3.3. Comparing Modules
    URR traverses the list of hashes of the top-level nodes that represent the module observed in the wild.
@@
    It looks up the hash of each node in the database and identifies every library that has the same top-level node. It considers any such library as a candidate library.
@@
    It appends the weight of the node for each candidate library to previous matches, if any. Therefore, the weights associated with a candidate library increases each time it includes a match.
@@ 3.4. Replacing Modules
URR replaces each marked privacy-harming module with a corresponding, benign replacement. It can perform this action in one of two ways; (1) URR can replace the module’s AST with an alternative, benign AST and then regenerate the script from from the modified AST; (2) Alternatively, URR can identify the string indices for the module within the textual representation of the script, and place the benging replacement between these indices. Regardless of the approach it adopts, URR ensures that access and use of the replacement does not break other parts of website functionality.
Replacements for targeted modules are manually created. Given the nuances of each target library and its eventual, mutated representation in webpack bundles, we describe the three categories of replacements to consider within each module.
  • Global Variables. If libraries expose global variables on the window object, these variables need to be made available. Additionally, the types of such variables need to be retained.
  • Exports. The replacement needs to ensure that values previously exported from the target module remain available. When the module is consumed by other modules, they need to be made accessible, even if replaced with a benign version.
  • Webpack-based Function Wrappers. When webpack wraps modules in functions, it passes specific parameters included in the function signature. Additionally, values exported by modules are appended as properties to webpack objects. Replacements need to ensure that these parameters and exports work seemlessly.
@@ 4. Evaluation
In this section we evaluate the effectiveness of URR in identifying specific libraries in bundled scripts captured in the wild. We select three libraries to target and describe the process of gathering their module representations for comparison. We then use URR to identify target libraries in scripts gathered from a crawl of popular websites. Next, we develop a deployment of URR as a Firefox extension and evaluate the performance overhead introduced by its interventions. We conclude by describing the creation and evaluation of benign replacements.
@@ 4.1. Evaluation Dataset
We first describe the selection of example libraries that we use to identify in the wild. We describe the offline process of gathering representations for these libraries which URR can then use to identify libraries in the wild.
@@ 4.1.1. Target Library Selection
We select three libraries for our evaluation —FingerprintJS (FingerprintJS, 2023a), Sentry (Sentry, 2023), and Prebid (Prebid, 2023). Each of these libraries are included in filter lists (EasyList Authors, 2023; EasyPrivacy Authors, 2023) used by popular content blocking tools (Hill and Rolls, 2023; eyeo GmbH, 2023). Additionally, all three libraries provide the option for developers to use their npm packages with bundled code (Fin, 2023b; Sen, 2023; Pre, 2023).
FingerprintJS (FingerprintJS, Inc, 2023) is a popular browser fingerprinting library that tracks and identifies users. When embedded within a website that a user visits, the library performs various client-side operations that query browser attributes, and stores a unique identifier for the user. Besides its use in user authentication (Lin et al., 2022) and ad fraud prevention (Fin, 2023a), the library de-anonymizes sensitive user activity across sites. Since the domains that the library uses are blocked by popular content blocking tools, FingerprintJS recommends that developers use their npm package or self-host a copy of their scripts (FingerprintJS, 2023b).
Sentry (Functional Software Inc., 2023) is an analytics tool that offers libraries for performance and error monitoring. The library helps developers gathers details about user interaction, DOM events, console logs, and network calls. The library lets developers decide how they use the information gathered from users and offers multiple integrations, including with multiple third-party analytics services (Sentry, 2023a). Since the CDNs that host by Sentry code are blocked by popular content blockers, the library recommends that developers get around this restriction by bundling Sentry’s npm package into their app (Sentry, 2023b).
Prebid (Prebid.org Inc, 2023) is a popular advertising library that developers can use to add header bidding to their application. The library integrates with numerous advertising, analytics, and user tracking libraries, providing support for bidding on targeted advertisements. Similar to previously discussed libraries, Prebid is included within filter lists and content blocking tools. The library is open source and is available to be added to bundles with an npm package. Prebid provides specific instructions on integration with webpack (Prebid, 2023; Babel, 2023). The library needs to be compiled with Babel and makes use of specific plugins when loaded with webpack. It provides specific instructions for webpack configurations, which gives us insight into its use when integrated by bundles in the wild.
Table 2. An overview of of the build options used to gather representations for the target library.
@@
Web applications are developed in a large variety of environments, frameworks, and configurations, before they are compiled and shipped to production. As a result, when the same npm package is bundled within different applications, its modules will generate a wide range of AST structures. These variations make it difficult to gather a definitive AST representation for the package that will return perfect matches when compared against scripts found in the wild. However, the AST of the module corresponding to each library is complex and hence, unique. Our intuition is that the AST of a given library will be closer to differently configured ASTs of the same library than to ASTs of other libraries.
@@
    Module Systems. Depending on the target npm package and the development evironment of the web application, the library may be consumed as a CommonJS module or an EcmaScript module. We gather configurations for both module types.
@@
    Webpack Optimizations. When bundling code from multiple files, webpack keeps track of imported and exported values with the help of a dependency graph. Depending on the module system, the output environment, and developer’s configurations, it provides options to automatically identify and discard unused parts of code and to reduce the size of the final output. The use of these options can change how the module appears in the output bundle.
@@
    Minifier Options. Once webpack creates a bundle output, it uses Terser (Terser, 2023) to minify the code. Developers can provide additional options to ensure that the output works with their intended production environment, ensuring that their applications work in all supported environments. We consider a list of compression options that alter the syntax of the module in the bundled output.
@@
    Target Module. When building a bundle, webpack provides the option to gather relevant info about individual modules (webpack, 2023). We use this information to identify the file, name, and position of the module corresponding to the target library. Thereafter, we gather an AST representation of the output bundle and extract the sub-tree corresponding to the target module.
@@
In addition to the three target libraries, we gathered a list of 10k popular npm packages based on their download counts (npm/registry, 2023; Kashcha, 2023; F, 2023). From this list, we collected a random sample of 10 versions each of 1k libraries. For each sampled library version, we generated 100 webpack bundles with each with a different, random build option. We extracted and processed their modules and gathered a weighted list of hashes corresponding to the top-level nodes of their ASTs.
4.2. Replacements
Prior work has extensively studied and determined the creation and use of non-breaking, benign replacements for privacy-harming code (Smith et al., 2021). We follow similar principles and extend example replacements used by uBlock Origin (uBl, 2023a).
4.2.1. Creating Benign Versions of Target Libraries.
For each of the three libraries, we considered associated source code and documentation to create equivalent benign versions. We briefly describe them below.
FingerprintJS, when loaded as a library from its npm package, provides a function, load(), which returns a Promise that resolves to a computed visitorId. While the computation includes privacy-harming code, the benign replacement that we developed only returns a randomly generateed visitorId. We used a similar approach to the existing replacment included within uBlock origin (uBl, 2023b), but modified the same to a Node.js module.
Sentry provides a browser library as an entry point npm package (Sen, 2023) to include its functionality within web applications. The library further includes other related packages and integrations that vary across developer and application use cases. We used the official documentation as a guideline and developed replacements for exported values, included those imported from other Sentry libraries. For each class that the library exports, we developed an equivalent bengign class with a constructor that accepts any argument it is passed. Similarly, we developed benign function replacements, with no return value, for each exposed function. For functions that return promises, we resolve into an empty promise. Finally, since the library additionally makes functionality available globally, via the window object, we also make the benign version of the library globally available.
Prebid provides a similar npm package to sentry, i.e., its main functionality is included from an entry point module, while developers can make use of additional integrations. We focus on the entry point module, i.e., pbjs, since all other functionality is accessed from this module. We follow a similar approach, i.e., creating benign replacements for each function and variable made available on the entry point module, the Prebid API for publishers (i.e., web application developers), and ensure that any access to the benign version is considered valid and does not throw an error when accessed.
4.2.2. Gathering Bundled Versions.
For each replacment module, we created a barebones application that accesses the library functionality. The barebones application helps evaluate each replacement, ensuring that the application remains functional. We package the resulting application and gathered the webpack module corresponding to the benign replacement, for use in URR’s deployment.
@@ 4.3. Framework Effectiveness
In this section, we describe our evaluation of URR’s effectiveness in identifiying target libraries within scripts captured in the wild.
@@
We used a puppeteer-based crawler that visited sites from the the Tranco list (Le Pochat et al., 2019) and gathered JavaScript resources through network interception. We previously used this dataset to create a bundle parser, described in Section 3.1. The dataset comprises 30,930 scripts from 1,063 sites, of which 11,995 scripts are bundled resources.
@@
We observed the tradeoff between precision and recall for the matches returned for all three target libraries. We looked for the lowest threshold match percentage for an AST for which all three libraries observed 100% precision, i.e., any AST that returned a positive match above this threshold was correctly annotated. We therefore arrived at a common match threshold value of 8%. Despite observing a small drop in recall for Sentry, we prioritized precision to ensure that URR never blocks benign modules even if it misses some privacy-harming scripts.
@@
Next, we performed a larger crawl of the web, gathering scripts from the top 10K sites in the Tranco list. We used URR to evaluate these scripts and employed the ¿=8% match threshold discussed earlier. We detected FingerprintJS on 205 sites, Sentry on 213 sites, and Prebid on 325 sites. Overall, URR identified at least one of the three target libraries on 7% ($n=697$) of the top 10K sites. Figure 3 presents an overview of our observations.

### v3

@@ 3.1.1. Gathering JavaScript resources
To create a valid parser for bundles in the wild, we gathered examples of resources loaded on popular websites. We developed a puppeteer-based crawler that, upon visiting a page, intercepts network requests and stores a copy of observed script responses. We visited domains from the Tranco list (Le Pochat et al., 2019) and gathered 30,930 scripts from 1,063 sites (1K crawl).
@@ 3.1.2. Generating ASTs
For each of the gathered scripts, URR uses acorn, a community-developed, open-source JavaScript parser to generate an AST [^10] https://github.com/acornjs/acorn. The generated AST complies with the ESTree Spec to ensure a consistent, standardized representation that can be reproduced by alternative implementations [^11] https://github.com/estree/estree.
@@
We began with an understanding of the general structure of the output of a webpack bundle (see Listing 1). We developed a script that parses an AST and looks for a module array or object, i.e., an array or object that comprises functions. We used this initial logic and adopted a code optimization methodology based on ground truth gathered through expert manual annotation. We describe the process below.
@@
    We manually evaluated the plaintext script and the AST of each sampled resource. In doing so, we checked for the presence of webpack-specific code patterns. We annotated each resource with a boolean value indicating whether we determined the script as a webpack bundle (Annotation 1). Note that these manual checks (e.g.,keywords) were only used for annotation and not incorporated into the automated approach. The manual checks included:
@@
    We used the refined version of our code to annotate the same sample and noted any improvements. This refined code version is the initial version of the subsequent phase against new samples.
@@
We used the resulting logic to annotate all scripts gathered in our 1K crawl. We annotated a total of 11,995 scripts as webpack bundles. Finally, we performed a spot check of the generated annotation. We randomly sampled and manually annotated 300 previously unsampled scripts, of which we identified 99 scripts as webpack bundles. Upon comparing and verifying the code-assigned annotation, we observed a 97% precision rate in identifying webpack bundles and a 99.8% precision rate in gathering component modules.
@@ 3.2. Processing Modules
Each module gathered from the previous phase comprises a sub-tree, i.e., a partial AST of the larger AST representation of the script. The module’s AST contains information about the structure of the underlying code and the associated names of variables, functions, and properties included within the script. Script attributes like variable names are extremely volatile and have limited use in reliably identifying target modules. URR, therefore, only considers (1) the structure of the AST, i.e., the parent-child relationships between the nodes that comprise the AST, and (2) the type of each AST node, i.e., attributes which comply with the ESTree Spec, and are hence limited to a deterministic set of values.
To distill and represent only these specific attributes from a module’s AST, we adopt a version of a cryptographic concept used in integrity verification: Merkle trees (Merkle, 1987). This approach presents a novel part of our methodology. For example, consider a list of data blocks that need to be verified (e.g., transactions in a blockchain). We represent this list of blocks in a tree structure. Consider a tree where each leaf node represents a piece of data, and each non-leaf node is a cryptographic hash of its child nodes. These hashes propagate upwards, converging into a singular root hash known as the Merkle root. Any alteration in the foundational data triggers a modification in the root hash, instantly signaling tampering or modifications. However, if the root hash is verified, all of its children are also verified.
@@
    For every child node, it recursively calls the function to gather the child node’s hash. The hash of each child, in turn, results from the hashes of its children.
@@
    To generate a hashed representation of the root node, URR concatenates the type of the node to the sum of hashes of its children and hashes the resulting string.
@@
This processed representation now comprises an alternative tree representation with equal depth to its original counterpart. However, tree-based comparisons are complex and add a large performance overhead. Recall that while processing an AST, top-level nodes contain information about underlying children. Therefore, comparing a limiting comparison to a few high-level nodes of the processed module can provide insight into the similarity of the module’s AST. To this end, URR gathers a list of top-level nodes. In other words, noting that webpack wraps each module within a function, we consider the computed hashes of each statement within the function. Say, for example, when bundled within webpack, the module comprises three statements: one variable declaration statement, one function definition (which further contains multiple nested statements), and one export statement. URR will process this module and gather hashes for three top-level nodes – owing to the Merkle tree-based approach, each of the hashes will be computed from the top-level statements’ children.
Additionally, URR associates each entry in this list with a weight that represents the number of child nodes it represents, i.e., it weighs a node representing a complex function higher than a node representing a simple variable declaration. Consider the same example as above. Here, the initial declaration statement and the export statement will represent fewer nodes than the function definition (which has children within nested statements). As a result, the function declaration statement will have a higher weight, corresponding to the ratio of nodes it represents to the total nodes in the module’s overall AST.
@@ 3.3. Comparing Modules
    URR traverses the list of hashes of the top-level nodes representing the module observed in the wild.
@@
    It looks up the hash of each node in the database and identifies every library with the same top-level node. It considers any such library as a candidate library.
@@
    It appends the node’s weight for each candidate library to previous matches, if any. Therefore, the weights associated with a candidate library increase each time it includes a match.
@@ 3.4. Replacing Modules
URR replaces each marked privacy-harming module with a corresponding, benign replacement. It can perform this action in one of two ways: (1) URR can replace the module’s AST with an alternative, benign AST and then regenerate the script from the modified AST; (2) Alternatively, URR can identify the string indices for the module within the textual representation of the script, and place the benign replacement between these indices. Regardless of its approach, URR ensures that access and use of the replacement does not break other parts of website functionality.
Replacements for targeted modules are manually created. Given the nuances of each target library and its eventual mutated representation in webpack bundles, we describe the three categories of replacements to consider within each module.
  • Global Variables. If libraries expose global variables on the window object, these variables must be made available. Additionally, the types of such variables need to be retained.
  • Exports. The replacement must ensure that values previously exported from the target module remain available. When other modules consume the module, it must be made accessible, even if replaced with a benign version.
  • Webpack-based Function Wrappers. When webpack wraps modules in functions, it passes specific parameters included in the function signature. Additionally, values exported by modules are appended as properties to webpack objects. Replacements need to ensure that these parameters and exports work seamlessly.
@@ 4. Evaluation
In this section, we evaluate the effectiveness of URR in identifying specific libraries in bundled scripts captured in the wild. We select three libraries to target and describe the process of gathering their module representations for comparison. We then use URR to identify target libraries in scripts gathered from a crawl of popular websites. Next, we develop a deployment of URR as a Firefox extension and evaluate the performance overhead introduced by its interventions. We conclude by describing the creation and evaluation of benign replacements.
@@ 4.1. Evaluation Dataset
We first describe the selection of example libraries that we use to identify in the wild. We describe the offline process of gathering representations for these libraries, which URR can then use to identify libraries in the wild.
@@ 4.1.1. Target Library Selection
We select three libraries for our evaluation —FingerprintJS [^12] https://github.com/fingerprintjs/fingerprintjs, Sentry [^13] https://github.com/getsentry/sentry-javascript, and Prebid [^14] https://github.com/prebid/Prebid.js. Each of these libraries is included in filter lists [^15] https://easylist.to/ used by popular content blocking tools [^16] https://github.com/gorhill/uBlock [^17] https://adblockplus.org/. Additionally, all three libraries provide the option for developers to use their npm packages with bundled code.
FingerprintJS [^18] https://fingerprint.com/ is a popular browser fingerprinting library that tracks and identifies users. When embedded within a website that a user visits, the library performs various client-side operations that query browser attributes and stores a unique identifier for the user. Besides its use in user authentication (Lin et al., 2022) and ad fraud prevention [^19] https://fingerprint.com/case-studies/, the library de-anonymizes sensitive user activity across sites. Since the domains that the library uses are blocked by popular content-blocking tools, FingerprintJS recommends that developers use their npm package or self-host a copy of their scripts [^20] https://github.com/fingerprintjs/fingerprintjs/blob/master/docs/evade_ad_blockers.md.
Sentry [^21] https://sentry.io/ is an analytics tool that offers libraries for performance and error monitoring. The library helps developers gather details about user interaction, DOM events, console logs, and network calls. The library lets developers decide how they use the information gathered from users and offers multiple integrations, including multiple third-party analytics services [^22] https://develop.sentry.dev/development/analytics/. Since the CDNs that host Sentry code are blocked by popular content blockers, the library recommends that developers get around this restriction by bundling Sentry’s npm package into their app [^23] https://docs.sentry.io/platforms/javascript/troubleshooting/#dealing-with-ad-blockers.
Prebid [^24] https://prebid.org/ is a popular advertising library that developers can use to add header bidding to their applications. The library integrates with numerous advertising, analytics, and user-tracking libraries, providing support for bidding on targeted advertisements. Similar to previously discussed libraries, Prebid is included within filter lists and content-blocking tools. The library is open source and is available to be added to bundles with an npm package. Prebid provides specific instructions on integration with webpack [^25] https://github.com/prebid/Prebid.js. The library must be compiled with Babel [^26] https://babeljs.io/, and specific plugins must be used when loaded with webpack. It provides specific instructions for webpack configurations, which gives us insight into its use when integrated by bundles in the wild.
Table 2. An overview of the build options used to gather representations for the target library.
@@
Web applications are developed in a large variety of environments, frameworks, and configurations before they are compiled and shipped to production. As a result, when the same npm package is bundled within different applications, its modules will generate a wide range of AST structures. These variations make it difficult to gather a definitive AST representation for the package that will return perfect matches when compared against scripts found in the wild. However, the AST of the module corresponding to each library is complex and, hence, unique. Our intuition is that the AST of a given library will be closer to differently configured ASTs of the same library than to ASTs of other libraries.
@@
    Module Systems. Depending on the target npm package and the development environment of the web application, the library may be consumed as a CommonJS module or an ECMAScript module. We gather configurations for both module types.
@@
    Webpack Optimizations. When bundling code from multiple files, webpack keeps track of imported and exported values with the help of a dependency graph. Depending on the module system, the output environment, and the developer’s configurations, it provides options to automatically identify and discard unused parts of code and to reduce the size of the final output. The use of these options can change how the module appears in the output bundle.
@@
    Minifier Options. Once webpack creates a bundle output, it uses Terser [^27] https://terser.org/ to minify the code. Developers can provide additional options to ensure that the output works with their intended production environment and that their applications work in all supported environments. We consider a list of compression options that alter the syntax of the module in the bundled output.
@@
    Target Module. When building a bundle, webpack provides the option to gather relevant info about individual modules [^28] https://webpack.js.org/api/stats/. We use this information to identify the file, name, and position of the module corresponding to the target library. Thereafter, we gather an AST representation of the output bundle and extract the sub-tree corresponding to the target module.
@@
In addition to the three target libraries, we gathered a list of 10k popular npm packages based on their download counts [^29] https://github.com/anvaka/npmrank. From this list, we collected a random sample of 10 versions of each of the 1k libraries. For each sampled library version, we generated 100 webpack bundles each with a different, random build option. We extracted and processed their modules and gathered a weighted list of hashes corresponding to the top-level nodes of their ASTs.
@@ 4.2. Framework Effectiveness
In this section, we describe our evaluation of URR’s effectiveness in identifying target libraries within scripts captured in the wild.
@@
We used a puppeteer-based crawler that visited sites from the Tranco list (Le Pochat et al., 2019) and gathered JavaScript resources through network interception. We previously used this dataset to create a bundle parser, described in Section 3.1. The dataset comprises 30,930 scripts from 1,063 sites, of which 11,995 scripts are bundled resources.
@@
We observed the tradeoff between precision and recall for the matches returned for all three target libraries. We looked for the lowest threshold match percentage for an AST for which all three libraries observed 100% precision, i.e., any AST that returned a positive match above this threshold was correctly annotated. We, therefore, arrived at a common match threshold value of 8%. Despite observing a small drop in recall for Sentry, we prioritized precision to ensure that URR never blocks benign modules even if it misses some privacy-harming scripts.
@@
Next, we performed a larger crawl of the web, gathering scripts from the top 10K sites in the Tranco list. We used URR to evaluate these scripts and employed the ¿=8% match threshold discussed earlier. We detected FingerprintJS on 205 sites, Sentry on 213 sites, and Prebid on 325 sites. Overall, URR identified at least one of the three target libraries on 7% ($n=697$) of the top 10K sites. Figure 3 presents an overview of our observations. We note that the larger crawl has fewer matches than the Top-1K sites, which can possibly be explained by existing work, which states that (a) less popular websites are less complex (Ruth et al., 2022), and (b) popular sites are more likely to engage in privacy-harming behaviors (Papadogiannakis et al., 2021).